OpenStudy (anonymous):

brandon recieved a book for his birthday and read 63 pages in 3hours later he read 84 pages in 4hours does his reading follow a direction variation explain

OpenStudy (anonymous):

\[\frac{ 63 }{ 3 } = \frac{ 21 }{ 1 }\] \[\frac{ 84 }{ 4 } =\frac{ 21 }{ 1 }\]They both simplify to the same ratio.